OpenStudy (anonymous):

Sam is building a model of an antique car. The scale of his model to the actual car is 1:10. His model is 18 1/2 inches long. How long is the actual car? Plz help me

OpenStudy (anonymous):

make a ratio 1/10 = 18.5/x x = 18.5(10) = 185 inches long